# Deep Reinforcement Learning for Wireless Scheduling in Distributed Networked Control

### v4

@@ III Stability Condition
Before turning to designing scheduling policies, it is critical to elucidate conditions that the WNCS needs to satisfy which ensure that there exists at least one stationary and deterministic scheduling policy that can stabilize all plants using the available network resources. Note that a policy $\pi$ is a function mapping from a state to an action. A deterministic policy means that the function gives the same action when the input state is fixed. A stationary policy means the function $\pi$ is time invariant, i.e., $\pi_{k}=\pi,\forall k$. In other words, the action only depends on the state, not the time [28]. We adopt a very commonly considered stochastic stability condition as below.
@@
Intuitively, the stability condition of the WNCS should depend on both the (open-loop) unstable plant systems (i.e., those where $\rho(\mathbf{A}_{i})\geq 1$) and the $M$-frequency communication system parameters. For the tractability of sufficient stability condition analysis (i.e., to prove the existence of a stabilizing policy), we focus on policy class that groups the unstable plants into $M$ disjoint sets $\mathcal{F}_{1},\dots,\mathcal{F}_{M}$ and allocates them to the $M$ frequencies, accordingly. Let ¯$\bar{\mathcal{F}}\triangleq\{i:\rho(\mathbf{A}_{i})\geq 1,i\in\{1,\dots,N\}\}$ denote the index set of all unstable plants. We have $\mathcal{F}_{1}\cup\dots\cup\mathcal{F}_{M}=\bar{\mathcal{F}}\subseteq\{1,\dots,N\}$ and $\mathcal{F}_{i}\cap\mathcal{F}_{j}=\emptyset,\forall i\neq j$. Then, we present the stability condition below which takes into account all potential allocations $(\mathcal{F}_{1},\dots,\mathcal{F}_{M})$.
@@
$\kappa\triangleq\min_{(\mathcal{F}_{1},\dots,\mathcal{F}_{M})}\max_{m=1,\dots,M,\mathcal{F}_{m}\neq\emptyset}\rho^{\max}_{m}\bar{\xi}^{\max}_{m}<1.$ (21)
@@
We construct a multi-frequency persistent scheduling policy: the unstable plants are grouped into $M$ sets $\mathcal{F}_{1},\dots,\mathcal{F}_{M}$, corresponding to the $M$ frequencies. In each frequency, the controller schedules the uplink of the first plant, say plant $i$, persistently until success. It then persistently schedules the downlink until success, and waits for $v_{i}-1$ steps for applying all the control commands in the actuator’s buffer. It then schedules the uplink of the second plant and so on and so forth. This procedure is repeated ad-infinitum. The reason we choose such a policy is because of its tractability and the tightness of the sufficient stability condition that we will derive. The detailed proof is included in [42]. ∎
@@
Theorem 1 captures the stabilizability of the WNCS scheduling problem in terms of both the dynamic system parameters, $\mathbf{A}_{i},\forall i\in\bar{\mathcal{F}}$, and the wireless channel conditions, i.e., $\{\bar{\xi}^{s}_{m,i},\bar{\xi}^{c}_{m,i}\},i\in\bar{\mathcal{F}},m=1,\dots,M$. Once (21) holds, there is at least one stationary and deterministic policy that stabilized all plants of the WNCS. If the channel quality of different links does not differ much at the same frequency, then the sufficient stabilizability condition is tight. To the best of our knowledge, this is the first stabilizability condition established for $N$-plant-$M$-frequency WNCS with uplink and downlink scheduling in the literature.
@@ IV-A MDP State Definition
to record the estimation quality at $k$ and at the $v_{i}$ successful control transmissions, where $\tau_{i,k}^{0}\triangleq\tau_{i,k}$ was defined above (11). Similarly, we define
@@ IV-C Resulting MDP
The computation complexity of relative value iteration algorithm is given as $\mathcal{O}(|\mathcal{S}_{L}||\mathcal{A}|^{2}K)$ [43], where $K$ is the number of iteration steps, and the state space and action space sizes are $|\mathcal{S}|=L^{2\sum_{i=1}^{N}v_{i}+2N}$ and $|\mathcal{A}|=\sum_{m=0}^{M}\mathsf{C}^{m}_{M}\mathsf{P}^{m}_{2N}$, respectively. However, the sizes of both the state and action spaces are considerably large even for relatively small $N$ and $M$, leading to numerical difficulties in finding a solution. In the literature of WNCS, even for the (simpler) $N$-plant-$M$-channel remote estimation system, only the $M=1$ case has been found to have a numerical solution [9]. To tackle the challenge for larger scale WNCS deployment, we will use DRL methods exploiting deep neural networks for function approximation in the following.
@@ V WNCS Scheduling with Deep Reinforcement Learning
Building on the MDP framework, DRL is widely applied in solving decision making problems with pre-designed state space, action space, per-step reward (cost) function and discount factor for achieving the maximum long-term reward. The main difference is that DRL does not exploit the state transition probability as required by MDP, but records and utilizes many sampled data sequences, including the current state $\mathbf{s}$, action $\mathbf{a}$, reward $r$ and next state $\mathbf{s}^{\prime}$, to train deep neural networks for generating the optimal policy [44]. To find deterministic policies [^5] The present work focuses on deterministic scheduling policies as it has been proved that the optimal policy is deterministic for unconstrained MDP problems [28]., the most widely considered DRL algorithms are DQN [45], DDPG [46], and TD3 [47]. In this study, we applied three algorithms to address the scheduling problem. DQN, which is the focus of our detailed algorithmic presentation, is the foundational algorithm in the field of deep reinforcement learning, primarily designed for environments with discrete action spaces. DDPG extends the ideas from DQN by adapting them to continuous action spaces using policy gradient methods. TD3, further building on DDPG, introduces mprovements in terms of stability and performance. DQN serves as a foundation for understanding the basic principles that are also applicable to the more complex algorithms like DDPG and TD3, and is well-suited for our problem (Section IV-C) with the discrete and finite action space $\mathcal{A}$. In this regard, we choose to describe the DQN algorithm in detailed technical depth to avoid redundancy. Detailed pseudocode for DDPG and TD3, while omitted from the main text to conserve space, is available in [46, 47] for interested readers. Although DDPG and TD3 are more typically applied to continuous action spaces, they can be adapted for discrete actions through the action embedding method introduced in this section.
@@
In general, we consider an episodic training scenario, where the system is operating in a simulated environment. Hence, one can learn and gather samples over multiple episodes to train a policy. Then, after verifying the properties of the final policy, one can “safely” deploy it [46]. We note that if the DRL policy does lead to instability (i.e., an unbounded expected cost), one needs to adjust the hyperparameters of the deep neural network (e.g., the initial network parameters and the number of layers and neurons) and start retraining. Such readjusting of hyperparameters of a deep learning agent is commonly adopted in practice [45].
@@ V-A Deep Q-Learning Approach
Solving the optimal $Q$ function is the key to find the optimal policy, but is computationally intractable by conventional methods as discussed in Remark 4. In contrast, deep Q-learning methods approximate $Q^{\star}(\mathbf{s,a})$ by a function $Q(\mathbf{s,a};\theta)$ parameterized by a set of neural network parameters, $\theta$, (including both weights and biases), and then learns $\theta$ to minimize the difference between the left- and right-hand sides of (46[48]. Deep Q-learning can be easily implemented by the most well-known machine learning framework TensorFlow with experience replay buffer, $\epsilon$-greedy exploration and mini-batch sampling techniques [49]. Building on this, the approach for solving problem (43) is given as Algorithm 1.
1:Initialize experience replay buffer $\mathcal{B}$ to capacity $K$
2:Initialize multi-layer fully connected neural network $Q$ with vector input $\mathbf{s}$, $|\mathcal{A}|$ outputs $\{Q(\mathbf{s},\mathbf{a}_{1};\theta_{0}),\dots,Q(\mathbf{s},\mathbf{a}_{|\mathcal{A}|};\theta_{0})\}$ and random parameter set $\theta_{0}$
3:for episode = $1,\cdots,E$ do
4:     Randomly initialize $\mathbf{s}_{0}$
5:     for $t=0,1,\cdots,T$ do
6:         With probability $\epsilon$ select a random action $\mathbf{a}_{t}$, otherwise select $\mathbf{a}_{t}=\arg\max_{\mathbf{a}\in\mathcal{A}}Q(\mathbf{s}_{t},\mathbf{a};\theta_{t})$
7:         Execute $\mathbf{a}_{t}$ , and obtain $r_{t}$ and $\mathbf{s}_{t+1}$
8:         Store $(\mathbf{s}_{t},\mathbf{a}_{t},r_{t},\mathbf{s}_{t+1})$ in $\mathcal{B}$
9:         Sample random mini-batch of $l$ transitions $(\mathbf{s}_{t},\mathbf{a}_{t},r_{t},\mathbf{s}_{t+1})$ from $\mathcal{B}$ as $\tilde{\mathcal{B}}$
10:         Set $z_{j}=r_{j}+\vartheta\max_{\mathbf{a}^{\prime}\in\mathcal{A}}\hat{Q}(\mathbf{s}_{j+1},\mathbf{a}^{\prime};\theta_{t})$ for each sample in $\tilde{\mathcal{B}}$
11:         Perform a mini-batch gradient descend step to minimize the Bellman error, i.e., $\min_{\hat{\theta}}\sum_{\tilde{\mathcal{B}}}(z_{j}-Q(\mathbf{s}_{j},\mathbf{a}_{j};\hat{\theta}))^{2}$
12:         Update $\theta_{t+1}=\hat{\theta}$
13:     end for
14:end for
Algorithm 1 Deep Q-learning for transmission scheduling in WNCS
@@ V-B Deep Q-Learning with Reduced Action Space
In practice, the training of deep Q-network (DQN) converges and the trained DQN performs well for simple decision-making problems with a bounded reward function, small input state dimension and small action space, see e.g., [49], where the state input is a length-$4$ vector and there are only $2$ possible actions. However, for the problem of interest in the current work, the reward function is unbounded (due to the potential of consecutive packet dropouts), the state $\mathbf{s}$ is of high dimension, and action-space $|\mathcal{A}|$ is large. Hence, the convergence of the DQN training is not guaranteed for all hyper-parameters, which include the initialization of $\theta$, the number of neural network layers, the number of neurons per layer, and the reply buffer and mini-batch sizes. [^6] Note that when $N=3$, $M=3$ and $v_{1}=\dots=v_{N}=2$, the length of the state is $18$ and the action space size is $229$ based on Section IV-C. Even if the convergence is achieved, due to the complexity of the problem, it may take very long training time and only converge to a local optimal parameter set $\theta$ (which is also affected by the choice of hyper-parameters), leading to a worse performance than some conventional scheduling policies. Whilst the hyper-parameters can in principle be chosen to enhance performance, no appropriate tuning guidelines exist for the problem at hand.
@@
The new scheduling action at the $M$ frequencies is denoted as $\breve{\mathbf{a}}_{k}\triangleq[\breve{a}_{1,k},\dots,\breve{a}_{M,k}]$, where $\breve{a}_{m,k}\in\{0,\dots,N\},\;\forall m\in\{1,\dots,M\}$, and $\breve{a}_{m,k}\neq\breve{a}_{m^{\prime},k}$ if $\breve{a}_{m,k}\breve{a}_{m^{\prime},k}\neq 0$. If $\breve{a}_{m,k}=i$ and $s^{\text{link}}_{i,k}=1$, then the uplink of plant $i$ is scheduled on frequency $m$ at time $k$. Compared with the original action $\mathbf{a}_{k}$ in Section IV-C, the size of the new action space is reduced significantly to $|\tilde{\mathcal{A}}|=\sum_{m=0}^{M}\mathsf{C}^{m}_{M}\mathsf{P}^{m}_{N}$. For example, when $N=3$, $M=3$ and $v_{1}=v_{2}=v_{3}=2$, the action space size is reduced from $229$ to $34$. By replacing $\mathbf{s}$ and $\mathbf{a}$ with $\breve{\mathbf{s}}$ and $\breve{\mathbf{a}}$, respectively, in Algorithm 1, we can apply the deep Q-learning method to find a policy with a reduced action space at a faster rate. It should be noted that the reduced action space still rapidly increases with the increment of $N$ and $M$. This makes it difficult for the DQN-based algorithm to solve the scheduling problem of larger-scale WNCS, such as $N=10$ and $M=10$ with a discrete action space of $234662231$.

### v5

@@ III Stability Condition
Before turning to designing scheduling policies, it is critical to elucidate conditions that the WNCS needs to satisfy to ensure that there exists at least one stationary and deterministic scheduling policy that can stabilize all plants using the available network resources. Note that a policy $\pi$ is a function mapping from a state to an action. A deterministic policy means that the function gives the same action when the input state is fixed. A stationary policy means the function $\pi$ is time invariant, i.e., $\pi_{k}=\pi,\forall k$. In other words, the action only depends on the state, not the time [31]. We adopt a very commonly considered stochastic stability condition as below.
@@
Definition 1 establishes mean‐square stability as the condition that the long‐term average of the expected squared state remains bounded. Our stability analysis is based on a stochastic framework that emphasizes average performance over time rather than focusing on instantaneous dropout events. In practice, there may be brief periods when $\gamma_{i,k}=0$, that is, when an actuator consecutively fails to receive new control commands. Although such short bursts can temporarily degrade performance, the overall system remains stable as long as these events occur infrequently enough. In other words, even if consecutive packet dropouts occur occasionally, the average packet success rate will ensure that the state does not grow unbounded. Thus, the stability condition guarantees that, over the long run, the system remains mean‐square stable under a properly designed scheduling protocol, rather than requiring stability at every single time instant. In the later analysis, we provide a specific scheduling policy with stability guarantees.
Intuitively, the stability condition of the WNCS should depend on both the (open-loop) unstable plant systems (i.e., those where $\rho(\mathbf{A}_{i})\geq 1$) and the $M$-frequency communication system parameters. For the tractability of sufficient stability condition analysis (i.e., to prove the existence of a stabilizing policy), we focus on a policy class that groups the unstable plants into $M$ disjoint sets $\mathcal{F}_{1},\dots,\mathcal{F}_{M}$ and allocates them to the $M$ frequencies, accordingly. Let ¯$\bar{\mathcal{F}}\triangleq\{i:\rho(\mathbf{A}_{i})\geq 1,i\in\{1,\dots,N\}\}$ denote the index set of all unstable plants. We have $\mathcal{F}_{1}\cup\dots\cup\mathcal{F}_{M}=\bar{\mathcal{F}}\subseteq\{1,\dots,N\}$ and $\mathcal{F}_{i}\cap\mathcal{F}_{j}=\emptyset,\forall i\neq j$. Then, we present the stability condition below which takes into account all potential allocations $(\mathcal{F}_{1},\dots,\mathcal{F}_{M})$.
@@
$\kappa\triangleq\min_{(\mathcal{F}_{1},\dots,\mathcal{F}_{M})}\max_{m=1,\dots,M,\mathcal{F}_{m}\neq\emptyset}\rho^{\max}_{m}\bar{\xi}^{\max}_{m}<1,$ (21)
where the operation $\min$ is taken over the set of all possible partitions $(\mathcal{F}_{1},\dots,\mathcal{F}_{M})$ of the set of unstable plants. For each such partition, we compute the maximum value over the $M$ frequencies, and then the minimum over all these partitions gives the most favorable (i.e., the smallest) worst-case product across the frequencies. The terms $\rho^{\max}_{m}$ and $\bar{\xi}^{\max}_{m}$ are computed with respect to the subset $\mathcal{F}_{m}$ of unstable plants.
@@
We construct a multi-frequency persistent scheduling policy: the unstable plants are grouped into $M$ sets $\mathcal{F}_{1},\dots,\mathcal{F}_{M}$, corresponding to the $M$ frequencies. In each frequency, the controller schedules the uplink of the first plant, say plant $i$, persistently until success. It then persistently schedules the downlink until success, and waits for $v_{i}-1$ steps for applying all the control commands in the actuator’s buffer. It then schedules the uplink of the second plant, and so forth. This procedure is repeated ad-infinitum. We choose such a policy because of its tractability and the tightness of the sufficient stability condition that we will derive. The detailed proof is included in [47]. ∎
Theorem 1 establishes that, once condition (21) is satisfied, there exists at least one stationary and deterministic scheduling protocol that guarantees mean-square stability of all plants. This does not imply that any arbitrary protocol will stabilize the system; rather, condition (21) ensures the existence of a suitable policy—such as the multi-frequency persistent scheduling approach we construct in the proof.
@@
Theorem 1 captures the stabilizability of the WNCS scheduling problem in terms of both the dynamic system parameters, $\mathbf{A}_{i},\forall i\in\bar{\mathcal{F}}$, and the wireless channel conditions, i.e., $\{\bar{\xi}^{s}_{m,i},\bar{\xi}^{c}_{m,i}\},i\in\bar{\mathcal{F}},m=1,\dots,M$. Once (21) holds, there is at least one stationary and deterministic policy that stabilizes all plants of the WNCS. If the channel quality of different links does not differ much at the same frequency, then the sufficient stabilizability condition is tight. To the best of our knowledge, this is the first stabilizability condition established for $N$-plant-$M$-frequency WNCS with uplink and downlink scheduling in the literature.
@@ IV-A MDP State Definition
to record the estimation quality at $k$ and at the $v_{i}$ successful control transmissions, where $\tau_{i,k}^{0}\triangleq\tau_{i,k}$ was defined above (11).
Similarly, we define
@@ IV-C Resulting MDP
The computation complexity of relative value iteration algorithm is given as $\mathcal{O}(|\mathcal{S}_{L}||\mathcal{A}|^{2}K)$ [48], where $K$ is the number of iteration steps, and the state space and action space sizes are $|\mathcal{S}|=L^{2\sum_{i=1}^{N}v_{i}+2N}$ and $|\mathcal{A}|=\sum_{m=0}^{M}\mathsf{C}^{m}_{M}\mathsf{P}^{m}_{2N}$, respectively. However, the sizes of both the state and action spaces are considerably large even for relatively small $N$ and $M$, leading to numerical difficulties in finding a solution. In the literature of WNCS, even for the (simpler) $N$-plant-$M$-channel remote estimation system, only the $M=1$ case has been found to have a numerical solution [9]. To tackle the challenge for larger scale WNCS deployment, we will use DRL methods exploiting DNNs for function approximation in the following.
@@ V WNCS Scheduling with Deep Reinforcement Learning
Building on the MDP framework, DRL is widely applied in solving decision making problems with pre-designed state space, action space, per-step reward (cost) function and discount factor for achieving the maximum long-term reward [49, 50]. The main difference is that DRL does not exploit the state transition probability as required by MDP, but records and utilizes many sampled data sequences, including the current state $\mathbf{s}$, action $\mathbf{a}$, reward $r$ and next state $\mathbf{s}^{\prime}$, to train DNNs for generating the optimal policy [51, 52]. To find deterministic policies [^4] The present work focuses on deterministic scheduling policies as it has been proved that the optimal policy is deterministic for unconstrained MDP problems [31]., the most widely considered DRL algorithms are DQN [53], DDPG [54], and TD3 [55]. In this study, we applied three algorithms to address the scheduling problem.
DQN, which is the focus of our detailed algorithmic presentation, is the foundational algorithm in the field of DRL, primarily designed for environments with discrete action spaces. DDPG extends the ideas from DQN by adapting them to continuous action spaces using policy gradient methods. TD3, further building on DDPG, introduces improvements in terms of stability and performance. DQN serves as a foundation for understanding the basic principles that are also applicable to the more complex algorithms like DDPG and TD3, and is well-suited for our problem (Section IV-C) with the discrete and finite action space $\mathcal{A}$. In this regard, we choose to describe the DQN algorithm in detailed technical depth to avoid redundancy. Detailed pseudocode for DDPG and TD3, while omitted from the main text to conserve space, is available in [54, 55] for interested readers.
Compared with standard DQN (see Section V-A), the primary novelty of our approach lies in how it tackles the rapid growth of the scheduling action space. First, we restrict each plant to either uplink or downlink mode, guided by the practical observation that sensor information must be received before control commands are updated (see Section V-B). Second, for scenarios involving many plants and frequencies, we embed these reduced actions into a continuous space by assigning a real-valued ‘priority score’ to each link (see Section V-C). This allows advanced DRL algorithms like DDPG or TD3 to be applied. Consequently, although we build upon the established DQN methodology in Section V-A, our innovations in action space reduction and action embedding enable us to solve large-scale scheduling problems more effectively than a naive DQN approach could.
@@
In general, we consider an episodic training scenario, where the system is operating in a simulated environment. Hence, one can learn and gather samples over multiple episodes to train a policy until convergence. Once training is completed offline, the properties of the final scheduling policy are verified so that one can “safely” deploy it at the central controller—rather than the sensors for real-time operation [54]. Online operation requires only a simple forward pass through the pre-trained neural network—a computationally lightweight procedure. The sensors merely provide state measurements and other necessary feedback, thereby avoiding any undue computational burden for sensors. We note that if the DRL policy does lead to instability (i.e., an unbounded expected cost), one needs to adjust the hyperparameters of the DNN (e.g., the initial network parameters and the number of layers and neurons) and start retraining. Such readjusting of hyperparameters of a deep learning agent is commonly adopted in practice [53].
@@ V-A Deep Q-Learning Approach
Solving the optimal $Q$ function is the key to finding the optimal policy, but it is computationally intractable by conventional methods as discussed in Remark 4. In contrast, deep Q-learning methods approximate $Q^{\star}(\mathbf{s,a})$ by a function $Q(\mathbf{s,a};\theta)$ parameterized by a set of neural network parameters, $\theta$, (including both weights and biases), and then learns $\theta$ to minimize the difference between the left- and right-hand sides of (46[56]. Deep Q-learning can be easily implemented by the most well-known machine learning framework, TensorFlow, with experience replay buffer, $\epsilon$-greedy exploration and mini-batch sampling techniques [57].
@@ V-B Deep Q-Learning with Reduced Action Space
In practice, the training of DQN converges and the trained DQN performs well for simple decision-making problems with a bounded reward function, small input state dimension and small action space, see e.g., [57], where the state input is a length-$4$ vector and there are only $2$ possible actions. However, for the problem of interest in the current work, the reward function is unbounded (due to the potential of consecutive packet dropouts), the state $\mathbf{s}$ is of high dimension, and action-space $|\mathcal{A}|$ is large. Hence, the convergence of the DQN training is not guaranteed for all hyper-parameters, which include the initialization of $\theta$, the number of neural network layers, the number of neurons per layer, and the replay buffer and mini-batch sizes. [^5] Note that when $N=3$, $M=3$ and $v_{1}=\dots=v_{N}=2$, the length of the state is $18$ and the action space size is $229$ based on Section IV-C. Even if the convergence is achieved, due to the complexity of the problem, it may take very long training time and only converge to a local optimal parameter set $\theta$ (which is also affected by the choice of hyper-parameters), leading to a worse performance than some conventional scheduling policies. Whilst the hyper-parameters can, in principle, be chosen to enhance performance, no appropriate tuning guidelines exist for the problem at hand.
@@
The new scheduling action at the $M$ frequencies is denoted as $\breve{\mathbf{a}}_{k}\triangleq[\breve{a}_{1,k},\dots,\breve{a}_{M,k}]$, where $\breve{a}_{m,k}\in\{0,\dots,N\},\;\forall m\in\{1,\dots,M\}$, and $\breve{a}_{m,k}\neq\breve{a}_{m^{\prime},k}$ if $\breve{a}_{m,k}\breve{a}_{m^{\prime},k}\neq 0$. If $\breve{a}_{m,k}=i$ and $s^{\text{link}}_{i,k}=1$, then the uplink of plant $i$ is scheduled on frequency $m$ at time $k$. Compared with the original action $\mathbf{a}_{k}$ in Section IV-C, the size of the new action space is reduced significantly to $|\tilde{\mathcal{A}}|=\sum_{m=0}^{M}\mathsf{C}^{m}_{M}\mathsf{P}^{m}_{N}$. For example, when $N=3$, $M=3$ and $v_{1}=v_{2}=v_{3}=2$, the action space size is reduced from $229$ to $34$. Building on this reduced action space, the approach for solving problem (43) is given as Algorithm 1. We note that the reduced action space still rapidly increases with the increment of $N$ and $M$. This makes it difficult for the DQN-based algorithm to solve the scheduling problem of larger-scale WNCS, such as $N=10$ and $M=10$ with a discrete action space of $234662231$.
1:Initialize experience replay buffer $\mathcal{B}$ to capacity $K$
2:Initialize multi-layer fully connected neural network $Q$ with vector input $\breve{\mathbf{s}}$, $|\tilde{\mathcal{A}}|$ outputs $\{Q(\breve{\mathbf{s}},\breve{\mathbf{a}}_{1};\theta_{0}),\dots,Q(\breve{\mathbf{s}},\breve{\mathbf{a}}_{|\mathcal{A}|};\theta_{0})\}$ and random parameter set $\theta_{0}$
3:for episode = $1,\cdots,E$ do
4:     Randomly initialize $\mathbf{s}_{0}$, and convert $\mathbf{s}_{0}$ to $\breve{\mathbf{s}}_{0}$
5:     for $t=0,1,\cdots,T$ do
6:         With probability $\epsilon$ select a random action $\breve{\mathbf{a}}_{t}$, otherwise select $\breve{\mathbf{a}}_{t}=\arg\max_{\breve{\mathbf{a}}\in\tilde{\mathcal{A}}}Q(\breve{\mathbf{s}}_{t},\breve{\mathbf{a}};\theta_{t})$
7:         Convert $\breve{\mathbf{a}}_{t}$ to $\mathbf{a}_{t}$
8:         Execute $\mathbf{a}_{t}$, and obtain $r_{t}$ and $\mathbf{s}_{t+1}$
9:         Convert $\mathbf{s}_{t+1}$ to $\breve{\mathbf{s}}_{t+1}$
10:         Store $(\breve{\mathbf{s}}_{t},\breve{\mathbf{a}}_{t},r_{t},\breve{\mathbf{s}}_{t+1})$ in $\mathcal{B}$
11:         Sample random mini-batch of $l$ transitions $(\breve{\mathbf{s}}_{t},\breve{\mathbf{a}}_{t},r_{t},\breve{\mathbf{s}}_{t+1})$ from $\mathcal{B}$ as $\tilde{\mathcal{B}}$
12:         Set $z_{j}=r_{j}+\vartheta\max_{\breve{\mathbf{a}}^{\prime}\in\tilde{\mathcal{A}}}\hat{Q}(\breve{\mathbf{s}}_{j+1},\breve{\mathbf{a}}^{\prime};\theta_{t})$ for each sample in $\tilde{\mathcal{B}}$
13:         Perform a mini-batch gradient descend step to minimize the Bellman error, i.e., $\min_{\hat{\theta}}\sum_{\tilde{\mathcal{B}}}(z_{j}-Q(\breve{\mathbf{s}}_{j},\breve{\mathbf{a}}_{j};\hat{\theta}))^{2}$
14:         Update $\theta_{t+1}=\hat{\theta}$
15:     end for
16:end for
Algorithm 1 Deep Q-learning with reduced action space for transmission scheduling in WNCS